# A generalized geometric mechanics theory for multi-curve-fold origami: vertex constrained universal configurations

### v1

@@ 2.3.2 Equilibrium equations for freely deformed systems
The variational principle indicates the Euler-Lagrange equations $f=0$ and $g=0$ along the crease, where
@@ 2.3.3 Equilibrium equations for systems with geometric constraints
In other cases, curved folding is conducted by applying geometric loads at both ends of the crease, such as fixing $\mathbf{r}_{c}(0)$ and $\mathbf{r}_{c}(S_{0})$ or giving closeness condition $\mathbf{r}_{c}(0)=\mathbf{r}_{c}(S_{0})$. Since no mechanical loads are applied on the panels, no edge-edge domain participates in the deformation, and the system is crease-dependent. The equilibrium state is also derived using the minimum total potential energy principle. The geometric constraints will introduce Lagrangian multipliers and thus make the Euler-Lagrange equations different from Eq. (31). However, many geometric constraints, including the displacement constraints, are not convenient to be expressed as functions of $\kappa$ and $\tau$. Instead, we variate $E$ by giving a virtual displacement $\delta\mathbf{r}_{c}$ of the crease. $\delta\kappa$ and $\delta\tau$ are then written as
@@
$f_{gc}$ $=f\kappa^{\prime}+2\kappa f^{\prime}+(\frac{-f\tau^{2}-f\kappa^{2}+f^{\prime\prime}}{\kappa})^{\prime}+g\tau^{\prime}+2\tau g^{\prime}+(\frac{2\tau(\frac{g^{\prime}}{\kappa})^{\prime}+\frac{g^{\prime}}{\kappa}\tau^{\prime}}{\kappa})^{\prime},$ (41)
$g_{gc}$ $=\frac{g^{\prime}}{\kappa}\tau^{2}-(g\kappa)^{\prime}-(\frac{g^{\prime}}{\kappa})^{\prime\prime}+f\tau^{\prime}+2\tau f^{\prime},$
@@ 3.2 Mechanics of multi-curve-fold origami
According to principles of variation, we obtained the Euler-Lagrange equations $f=0,g=0$ for structures without geometric constraints, where
@@ 4.1 Geometric constraints in single-vertex origami
The tangent vectors of folds intersecting at a vertex lead to extra geometric constraints of the folding angles at $S=0$, as shown in Fig. 5(b). We assume that generators cannot intersect at the vertex, otherwise the bending energy diverges. Under the assumption, $\langle\mathbf{t}_{i}(0),\mathbf{t}_{i+1}(0)\rangle$ stays $\alpha_{i}$ during isometric deformation. We define $\beta_{i}=\langle\mathbf{t}_{1}(0),\mathbf{t}_{i}(0)\rangle$, and according to the previous work on single-vertex origami with straight folds (Evans et al., 2015), $\phi_{i}(0)$, the folding angle of creases at the vertex, are expressed in forms of $\beta_{i}$ as
@@ 4.2 Near-field deformation
From $S_{1}=S_{3}$, we first derive $\kappa_{2}^{\prime}(S_{2})$. Given $\kappa_{1}(S_{1})$, $\kappa_{2}^{\prime\prime}(S_{2})$ can be derived by Eq. (56) and (57), and then $\tau_{2}^{\prime\prime}(S_{2})$ can be derived from $h_{\kappa 0}=0$. Inductively using Eq. (64), $\kappa_{2}^{(n)}$ and $\tau_{2}^{(n)}$ are derived. The curvature and torsion on the neighboring points are then expressed as
@@
We assume that $S_{3}>S_{1}$. As shown in Fig. 6(b), if the geometry on $(0,S_{2})$ of crease 2 is given, the geometry on $(0,S_{3})$ of crease 3 is determined according to Eq. (57) and (58). The geometry on $(0,S_{3})$ of odd folds then determines the geometry on $(0,S_{4})$ of even folds. The conduction will extend the geometry-determined domain, which continues until $\lim_{n\rightarrow\infty}(S_{n+2}-S_{n})=0$. Then we use Lemma 1 to extend the geometry-determined region, and the procedure continues until a generator reaches the boundary of the structure. ∎
@@
For folds with constant geodesic curvature, the coefficients of the series are easily derived via substituting Eq. (57) - (59) into (66).
@@ 4.4 Finite element modeling and numerical results of single-vertex curved origami
4.4.1 Finite element modeling of curved fold origami
@@ 5 Conclusion
In this work, a comprehensive framework for multi-curve-fold origami is established. This framework theoretically explains how geometry and elasticity determine the deformation of curved fold origami structures. As an example of multi-curve-fold origami with periodicity, the theory of single-vertex curved origami is established, unveiling a striking vertex-constrained universal equilibrium configuration. Our main contributions are summarized as follows:

### v2

@@ 2.3.2 Equilibrium equations for freely deformed systems
The variational principal indicates the Euler-Lagrange equations $f=0$ and $g=0$ along the crease, where
@@
The Euler-Lagrange equations in different domains are derived in this section, but solving the whole deformation of systems with different domains is challenging and needs further discussion. The domain distribution is generally unknown before solving the problem, thus all possible distributions and the related solutions should be considered. Furthermore, extra boundary conditions connecting different domains should be satisfied. A further discussion for single-curve-fold origami with different domains is presented in A, mainly focusing on the additional boundary conditions.
@@ 2.3.3 Equilibrium equations for systems with geometric constraints
In other cases, curved folding is conducted by applying geometric loads at both ends of the crease, such as fixing $\mathbf{r}_{c}(0)$ and $\mathbf{r}_{c}(S_{0})$ or giving closeness condition $\mathbf{r}_{c}(0)=\mathbf{r}_{c}(S_{0})$. Since no mechanical loads are applied on the panels, no edge-edge domain participates in the deformation, and the system is crease-dependent. The equilibrium state is also derived using the minimum total potential energy principal. The geometric constraints will introduce Lagrangian multipliers and thus make the Euler-Lagrange equations different from Eq. (31). However, many geometric constraints, including the displacement constraints, are not convenient to be expressed as functions of $\kappa$ and $\tau$. Instead, we variate $E$ by giving a virtual displacement $\delta\mathbf{r}_{c}$ of the crease. $\delta\kappa$ and $\delta\tau$ are then written as
@@
$f_{gc}$ $=f\kappa^{\prime}+2\kappa f^{\prime}+(\frac{-f\tau^{2}-f\kappa^{2}+f^{\prime\prime}}{\kappa})^{\prime}+g\tau^{\prime}+2\tau g^{\prime}+(\frac{2\tau(\frac{g^{\prime}}{\kappa})^{\prime}+\frac{g^{\prime}}{\kappa}\tau^{\prime}}{\kappa})^{\prime}=0,$ (41)
$g_{gc}$ $=\frac{g^{\prime}}{\kappa}\tau^{2}-(g\kappa)^{\prime}-(\frac{g^{\prime}}{\kappa})^{\prime\prime}+f\tau^{\prime}+2\tau f^{\prime}=0,$
@@ 3.2 Mechanics of multi-curve-fold origami
According to principals of variation, we obtained the Euler-Lagrange equations $f=0,g=0$ for structures without geometric constraints, where
@@ 4.1 Geometric constraints in single-vertex origami
The tangent vectors of folds intersecting at a vertex lead to extra geometric constraints of the folding angles at $S=0$, as shown in Fig. 5(b). We assume that generators cannot intersect at the vertex to form a conical point. Otherwise, $\kappa_{m}$ reaches infinity as discussed in Section 2.1, and the bending energy diverges. Under the assumption, $\langle\mathbf{t}_{i}(0),\mathbf{t}_{i+1}(0)\rangle$ stays $\alpha_{i}$ during isometric deformation. We define $\beta_{i}=\langle\mathbf{t}_{1}(0),\mathbf{t}_{i}(0)\rangle$, and according to the previous work on single-vertex origami with straight folds (Evans et al., 2015), $\phi_{i}(0)$, the folding angle of creases at the vertex, are expressed in forms of $\beta_{i}$ as
@@ 4.2 Near-field deformation
Given the initial values $\kappa_{2}(S_{2}),\tau_{2}(S_{2})$ and $\kappa_{1}(S_{1})$, the n-th order derivatives $\{\kappa_{2}^{(n)}(S_{2})\}$ and $\{\tau_{2}^{(n)}(S_{2})\}$ can be solved recursively according to Eq. (64) in the following procedure.
  1. 1.
    Since $S_{1}$ and $S_{3}$ are determined by $\kappa_{2}(S_{2})$, $\kappa_{2}^{\prime}(S_{2})$ and $\tau_{2}(S_{2})$ (see the equation below Eq. (64)), solving the assumption $S_{1}=S_{3}$ yields $\kappa_{2}^{\prime}(S_{2})$ as a function of $\kappa_{2}(S_{2})$ and $\tau_{2}(S_{2})$.
  2. 2.
    $\kappa_{1}(S_{1})$ is determined by $\kappa_{2}(S_{2}),\kappa_{2}^{\prime}(S_{2}),\kappa_{2}^{\prime\prime}(S_{2}),\tau_{2}(S_{2})$ and $\tau_{2}^{\prime}(S_{2})$ according to Eq. (46). $h_{\kappa 0}=0$ is also an equation for $\kappa_{2}^{\prime\prime}(S_{2})$ and $\tau_{2}^{\prime}(S_{2})$. Solving these two equations yields $\kappa_{2}^{\prime\prime}(S_{2})$ and $\tau_{2}^{\prime}(S_{2})$ as functions of the initial values.
  3. 3.
    The two equations $h_{\tau 0}=0$ and $h_{\kappa 1}=0$ give $\kappa_{2}^{\prime\prime\prime}(S_{2})$ and $\tau_{2}^{\prime\prime}(S_{2})$. Then the two equations $h_{\tau 1}=0$ and $h_{\kappa 2}=0$ give $\kappa_{2}^{(4)}(S_{2})$ and $\tau_{2}^{(3)}(S_{2})$.
  4. 4.
    Recursively, $\kappa_{2}^{(n+3)}(S_{2})$ and $\tau_{2}^{(n+2)}(S_{2})$ can be derived from $h_{\tau n}=0$ and $h_{\kappa(n+1)}=0$ by repeating the previous steps.
The curvature and torsion on the neighboring points are then expressed as
@@
We assume that $S_{3}>S_{1}$. As shown in Fig. 6(b), if the geometry on $(0,S_{2})$ of crease 2 is given, the geometry on $(0,S_{3})$ of crease 3 is determined according to Eq. (57) and (58). The geometry on $(0,S_{3})$ of odd folds then determines the geometry on $(0,S_{4})$ of even folds. The induction will extend the geometry-determined domain, which continues until $\lim_{n\rightarrow\infty}(S_{n+2}-S_{n})=0$. Then we use Lemma 1 to extend the geometry-determined region, and the procedure continues until a generator reaches the boundary of the structure. ∎
@@
For folds with constant geodesic curvature, the coefficients of the series are easily derived via substituting Eq. (57) - (59) into Eq. (66).
@@ 4.4 Finite element modeling and numerical results of single-vertex curved origami
4.4.1 Finite element modeling of curve-fold origami
@@ 5 Conclusion
In this work, a comprehensive framework for multi-curve-fold origami is established. This framework theoretically explains how geometry and elasticity determine the deformation of curve-fold origami structures. As an example of multi-curve-fold origami with periodicity, the theory of single-vertex curved origami is established, unveiling a striking vertex-constrained universal equilibrium configuration. Our main contributions are summarized as follows:
@@ Acknowledgement
Appendix A Further discussion on mechanics of single-curve-fold systems with different domains
In this appendix, we analyze the equilibrium equations and the boundary conditions for systems with different domains. While the domain classification is usually unknown before solving the equations, we should analyze all possible domain distributions and compare the related elastic energies. The distribution with the minimum energy is the real solution.
As stated in Section 2.3, the folding processes discussed in this paper will not introduce edge-edge domains. As analyzed in Section 2.2, the crease-crease domain only appears in systems with separate panels. We then give illustrations of possible generator distributions in a system that may have the crease-crease domain, the crease-edge domain and the plane domain, as shown in Fig. A.1. Let the starting point of the possible crease-crease domain $S_{1}^{*}$ be $S=0$. The value of $t(0)=(\kappa_{g}(0)\kappa_{n}^{\prime}(0)/\kappa^{2}(0)-\kappa_{n}(0)\kappa_{g}^{\prime}(0)/\kappa^{2}(0)+\tau(0))/\kappa_{n}(0)$ determines the type of the generator distribution. In the following sections, we give the total energy variations for both situations and then discuss the equilibrium equations and boundary conditions.
Figure A.1: Two possible generator distributions in a single-curve-fold system (panel $+$: red, panel --: blue). (a) If $t(0)<\cot{\theta^{*}}$, the system is covered with crease-edge domain and plane domain. (b) If $t(0)\geq\cot{\theta^{*}}$, the system is covered with crease-crease domain, crease-edge domain and plane domain. The black thick generator represents the boundary of the crease-crease domain.
A.1 Energy variations
$t(0)<\cot{\theta^{*}}$. In this situation, the system is covered with crease-edge domains and possible plane domains. Since the plane domain has no elastic energy, the elastic energy of the single-curve-fold system is
$E=\int_{0}^{S_{0}}\varepsilon(S){\textrm{d}}S,$ (A.1)
where $\varepsilon$ is defined in Eq. (29). Considering possible jumps in geometric variables caused by non-smooth edges and noncontinuous geodesic curvatures at $S_{jump}$, the energy variation in Eq. (30) is transformed into
$\delta E$ $=\int_{0}^{S_{0}}f\delta\kappa{\textrm{d}}S+\int_{0}^{S_{0}}g\delta\tau{\textrm{d}}S+(\partial_{\kappa^{\prime}}\varepsilon-(\partial_{\kappa^{\prime\prime}}\varepsilon)^{\prime})\delta\kappa\bigg{|}_{0}^{S_{0}}+\partial_{\kappa^{\prime\prime}}\varepsilon\delta\kappa^{\prime}\bigg{|}_{0}^{S_{0}}+\partial_{\tau^{\prime}}\varepsilon\delta\tau\bigg{|}_{0}^{S_{0}}+jump,$ (A.2)
where $f,g$ are defined in Eq. (31) and $jump$ is defined as
$jump=(\partial_{\kappa^{\prime}}\varepsilon-(\partial_{\kappa^{\prime\prime}}\varepsilon)^{\prime})\delta\kappa\bigg{|}_{S_{jump}-}^{S_{jump}+}+\partial_{\kappa^{\prime\prime}}\varepsilon\delta\kappa^{\prime}\bigg{|}_{S_{jump}-}^{S_{jump}+}+\partial_{\tau^{\prime}}\varepsilon\delta\tau\bigg{|}_{S_{jump}-}^{S_{jump}+}.$ (A.3)
$t(0)\geq\cot{\theta^{*}}$. In this situation, the system is covered with crease-crease domains, crease-edge domains and possible plane domains. the elastic energy of the system is
$E=\int_{0}^{S_{c}^{*}}\varepsilon^{*}(S){\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}\varepsilon(S){\textrm{d}}S,$ (A.4)
where $\varepsilon$ is defined in Eq. (29) and $\varepsilon^{*}$ is defined in Eq. (33). While $S_{2}$ is a moving boundary, the energy variation is
$\delta E$ $=\int_{0}^{S_{c}^{*}}f^{*}\delta\kappa{\textrm{d}}S+\int_{0}^{S_{c}^{*}}g^{*}\delta\tau{\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}f\delta\kappa{\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}g\delta\tau{\textrm{d}}S$ (A.5)
$+(\partial_{\kappa^{\prime}}\varepsilon^{*}-(\partial_{\kappa^{\prime\prime}}\varepsilon^{*})^{\prime})\delta\kappa\bigg{|}_{0}^{S_{c}^{*}}+\partial_{\kappa^{\prime\prime}}\varepsilon^{*}\delta\kappa^{\prime}\bigg{|}_{0}^{S_{c}^{*}}+\partial_{\tau^{\prime}}\varepsilon^{*}\delta\tau\bigg{|}_{0}^{S_{c}^{*}}$
$+(\partial_{\kappa^{\prime}}\varepsilon-(\partial_{\kappa^{\prime\prime}}\varepsilon)^{\prime})\delta\kappa\bigg{|}_{S_{2}^{*}+}^{S_{0}}+\partial_{\kappa^{\prime\prime}}\varepsilon\delta\kappa^{\prime}\bigg{|}_{S_{2}^{*}+}^{S_{0}}+\partial_{\tau^{\prime}}\varepsilon\delta\tau\bigg{|}_{S_{2}^{*}+}^{S_{0}}-\varepsilon(S_{2}^{*}+)\delta S_{2}^{*}+jump,$
where $f,g$ are defined in Eq. (31) and $f^{*},g^{*}$ are defined in Eq. (34).
As analyzed in Section 2.2, $\kappa_{n}(S_{c}^{*})=0,\kappa_{n}^{\prime}(S_{c}^{*})=0$ and $\tau(S_{c}^{*})=0$, leading to $\delta\kappa=0,\delta\kappa^{\prime}=0$ and $\delta\tau=0$ at $S_{c}^{*}$. Since a generator connects points $S=0$ and $S=S_{2}^{*}$, geometric correlations exist that $S_{2}^{*}$ is the function of $\kappa(0),\kappa^{\prime}(0),\tau(0)$ and $\kappa(S_{2}^{*}-)$ is the function of $\kappa(0),\kappa^{\prime}(0),\kappa^{\prime\prime}(0),\tau(0),\tau^{\prime}(0)$. To avoid energy singularity, $\kappa(S_{2}^{*}+)$ couples with $\kappa(S_{2}^{*}-)$ to keep the continuity of $\kappa/\kappa_{g}$. $\kappa(S_{2}^{*}+)$ is thus expressed as
$S_{2}^{*}=S_{2}^{*}(\kappa(0),\kappa^{\prime}(0),\tau(0)),$ (A.6)
$\kappa(S_{2}^{*}+)=\kappa(\kappa(0),\kappa^{\prime}(0),\kappa^{\prime\prime}(0),\tau(0),\tau^{\prime}(0)),$
leading to the variation
$\delta S_{2}^{*}=\frac{\partial S_{2}^{*}}{\partial\kappa(0)}\delta\kappa(0)+\frac{\partial S_{2}^{*}}{\partial\kappa^{\prime}(0)}\delta\kappa^{\prime}(0)+\frac{\partial S_{2}^{*}}{\partial\tau(0)}\delta\tau(0),$ (A.7)
$\delta\kappa(S_{2}^{*}+)=\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa(0)}\delta\kappa(0)+\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime}(0)}\delta\kappa^{\prime}(0)+\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime\prime}(0)}\delta\kappa^{\prime\prime}(0)+\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau(0)}\delta\tau(0)+\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau^{\prime}(0)}\delta\tau^{\prime}(0).$
Since the plane domain may occur at $S=S_{2}^{*}$ leading to discontinuous $\tau/\kappa_{g}$ and $\kappa^{\prime}/\kappa_{g}$, variations $\delta\kappa^{\prime}(S_{2}^{*}+)$ and $\delta\tau(S_{2}^{*}+)$ are free and irrelevant with $\delta\kappa^{\prime}(S_{2}^{*}-)$ and $\delta\tau(S_{2}^{*}-)$. The energy variation in Eq. (A.5) is then converted into
$\delta E$ $=\int_{0}^{S_{c}^{*}}f^{*}\delta\kappa{\textrm{d}}S+\int_{0}^{S_{c}^{*}}g^{*}\delta\tau{\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}f\delta\kappa{\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}g\delta\tau{\textrm{d}}S$ (A.8)
$-[(\partial_{\kappa^{\prime}}\varepsilon^{*}(0)-(\partial_{\kappa^{\prime\prime}}\varepsilon^{*}(0))^{\prime})+(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\kappa(0)}]\delta\kappa(0)$
$-[\partial_{\kappa^{\prime\prime}}\varepsilon^{*}(0)+(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime}(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\kappa^{\prime}(0)}]\delta\kappa^{\prime}(0)-[(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime\prime}(0)}]\delta\kappa^{\prime\prime}(0)$
$-[\partial_{\tau^{\prime}}\varepsilon^{*}(0)+(\partial_{\kappa^{\prime}}\varepsilon(S^{*}_{2}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S^{*}_{2}+))\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\tau(0)}]\delta\tau(0)-[(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau^{\prime}(0)}]\delta\tau^{\prime}(0)$
$+(\partial_{\kappa^{\prime}}\varepsilon(S_{0})-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{0}))^{\prime})\delta\kappa(S_{0})+\partial_{\kappa^{\prime\prime}}\varepsilon\delta\kappa^{\prime}\bigg{|}_{S_{2}^{*}+}^{S_{0}}+\partial_{\tau^{\prime}}\varepsilon\delta\tau\bigg{|}_{S_{2}^{*}+}^{S_{0}}+jump.$
A.2 Equilibrium equations and boundary conditions
From Eq. (A.5) and (A.8), we derive equilibrium equations and boundary conditions for both situations based on energy minimization $\delta E=0$.
$t(0)<\cot{\theta^{*}}$. For freely-deformed systems, $\delta\kappa$ and $\delta\tau$ are arbitrary. $\delta E=0$ thus leads to $f=0$ and $g=0$ satisfied along the crease $(0,S_{0})$, and the boundary conditions $\partial_{\kappa^{\prime}}\varepsilon-(\partial_{\kappa^{\prime\prime}}\varepsilon)^{\prime}=0$, $\partial_{\kappa^{\prime\prime}}\varepsilon=0$, $\partial_{\tau^{\prime}}\varepsilon=0$ on both ends $S=0,S=S_{0}$.
$jump=0$ introduces extra jump conditions at $S=S_{jump}$. Due to the continuity of $\kappa/\kappa_{g}$, $\delta\kappa(S_{jump}+)$ couples with $\delta\kappa(S_{jump}-)$, following $\delta\kappa(S_{jump}+)/\delta\kappa(S_{jump}-)=\kappa_{g}(S_{jump}+)/\kappa_{g}(S_{jump}-)$. In the contrast, $\delta\kappa^{\prime}(S_{jump}+)$ and $\delta\tau(S_{jump}+)$ are irrelevant with $\delta\kappa^{\prime}(S_{jump}-)$ and $\delta\tau(S_{jump}-)$. $jump=0$ thus leads to
$[\partial_{\kappa^{\prime}}\varepsilon(S_{jump}-)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{jump}-))^{\prime}]\kappa_{g}(S_{jump}-)-[\partial_{\kappa^{\prime}}\varepsilon(S_{jump}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{jump}+))^{\prime}]\kappa_{g}(S_{jump}+)=0,$ (A.9)
$\partial_{\kappa^{\prime\prime}}\varepsilon(S_{jump}+)=\partial_{\kappa^{\prime\prime}}\varepsilon(S_{jump}-)=0,$
$\partial_{\tau^{\prime}}\varepsilon(S_{jump}+)=\partial_{\tau^{\prime}}\varepsilon(S_{jump}-)=0.$
Satisfying the jump conditions in Eq. (A.9) may lead to jumps in $\tau$ and $\kappa^{\prime}$ at $S=S_{jumps}$, forming a plane domain.
$t(0)\geq\cot{\theta^{*}}$. For freely-deformed systems, $\delta\kappa$ and $\delta\tau$ are arbitrary. $\delta E=0$ thus leads to $f^{*}=0,g^{*}=0$ satisfied along $(0,S_{c}^{*})$ and $f=0,g=0$ satisfied along $(S_{2}^{*},S_{0})$. The boundary conditions are
$(\partial_{\kappa^{\prime}}\varepsilon^{*}(0)-(\partial_{\kappa^{\prime\prime}}\varepsilon^{*}(0))^{\prime})+(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\kappa(0)}=0$ (A.10)
$\partial_{\kappa^{\prime\prime}}\varepsilon^{*}(0)+(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime}(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\kappa^{\prime}(0)}]\delta\kappa^{\prime}(0)=0$
$(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\kappa^{\prime\prime}(0)}=(\partial_{\kappa^{\prime}}\varepsilon(S_{2}^{*}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{2}^{*}+))^{\prime})\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau^{\prime}(0)}=0$
$\partial_{\tau^{\prime}}\varepsilon^{*}(0)+(\partial_{\kappa^{\prime}}\varepsilon(S^{*}_{2}+)-(\partial_{\kappa^{\prime\prime}}\varepsilon(S^{*}_{2}+))\frac{\partial\kappa(S_{2}^{*}+)}{\partial\tau(0)}-\varepsilon(S_{2}^{*}+)\frac{\partial S_{2}^{*}}{\partial\tau(0)}=0$
$\partial_{\kappa^{\prime}}\varepsilon(S_{0})-(\partial_{\kappa^{\prime\prime}}\varepsilon(S_{0}))^{\prime}=0$
$\partial_{\kappa^{\prime\prime}}\varepsilon\delta\kappa^{\prime}\bigg{|}_{S_{2}^{*}+}^{S_{0}}=\partial_{\tau^{\prime}}\varepsilon\delta\tau\bigg{|}_{S_{2}^{*}+}^{S_{0}}=0$
Besides, the distributions of $\kappa,\tau$ should satisfy jump conditions in Eq. (A.9) at $S=S_{jump}$ and geometric constraints in Eq. (A.6). If the solutions exist for both situations, the system chooses the distribution of $\kappa$ and $\tau$ with the less elastic energy expressed in Eq. (A.1) and (A.4).
Besides the distributions illustrated in Fig. A.1, there are some special cases shown in Fig. A.2. If a plane domain shown in Fig. A.2(a) is formed, the crease-crease domain may not start from $S=0$ or $S=S_{0}$. Thus, the elastic energy is
$E=\int_{0}^{S_{1}^{*}}\varepsilon(S){\textrm{d}}S+\int_{S_{1}^{*}}^{S_{c}^{*}}\varepsilon^{*}(S){\textrm{d}}S+\int_{S_{2}^{*}}^{S_{0}}\varepsilon(S){\textrm{d}}S,$ (A.11)
where $S_{1}^{*}$ and $S_{2}^{*}$ are moving boundaries. If $\kappa_{g}$ changes its sign along the fold, the system may have more crease-crease domains, as shown in Fig. A.2(b). The elastic energy is
$E=\int_{0}^{S_{c}^{*}}\varepsilon^{*}(S){\textrm{d}}S+\int_{S_{2}^{*}}^{S_{1}^{**}}\varepsilon(S){\textrm{d}}S+\int_{S_{1}^{**}}^{S_{c}^{**}}\varepsilon^{*}(S){\textrm{d}}S,$ (A.12)
where $S_{2}^{*}$ and $S_{1}^{**}$ are moving boundaries. From (A.11) and (A.12), we can derive the equilibrium equations and boundary conditions for both cases, following the variational method given in this appendix.
Figure A.2: Two special cases of generator distributions in a single-curve-fold system (panel $+$: red, panel $-$: blue). (a) System with a plane domain formed between $(S_{1}^{*},S_{2}^{*})$. Thick generators represent the boundaries of the plane domain. (b) System with two crease-crease domains. $\kappa_{g}$ changes its sign along the fold. Thick generators represent boundaries of the crease-crease domains.
For systems with geometric constraints, the equilibrium equations and boundary conditions can be obtained by converting Eq. (35) into Eq. (A.5) and (A.8). Following Eq. (36) - (40), we can derive equilibrium equations satisfied in different domains (which we have studied in Section 2.3.3) and the boundary conditions.